A mass-on-a-spring system has a spring constant k = 360 N/m and a mass m = 0.53 kg. (a) If it is given an initial displacement of 0.27 m and then released, what is the initial potential energy of the oscillator? in J (b) What is the maximum kinetic energy of the oscillator? in J
Answer #1

a)  initial potential energy = 1/2KX^2

  initial potential energy = 1/2*360*(0.27)^2 = 13.12 J

b)  maximum kinetic energy of the oscillator = initial potential energy = 13.12 J ( as intial potential energy is max potential energy and as energy is conserved when PE = 0 , then whole PE will be converted to KE )